OpenStudy (anonymous):

If f(−4)=1 and g(x)=f(x−3), what point can you determine on the graph of g?

OpenStudy (danjs):

f(-4) = 1 is given , x=-4 and y=f(-4) = 1

OpenStudy (danjs):

you can determine g(x) = f(x-3) , when x=-1 here, because g(-1) = f(-1-3) = f(-4) = 1

OpenStudy (danjs):

g(-1) = 1

OpenStudy (anonymous):

Okay. That makes sense. Many thanks.

OpenStudy (danjs):

see how that worked?

OpenStudy (danjs):

The only point you know is f(-4), so to get g(x) you need the point where f(-4) is g(x) = f(x-3) to get f(-4) you have to let x=-1 g(-1) = f(-1-3) = f(-4) = 1